Find two consecutive integers such that the sum of 2 times the first integer and 4 times second integer is 46.
Answer by Paul(988)

Let the consecutive integers be x and x+1
Equation:
2(x)+4(x+1)=46
2x+4x+4=46
6x=42
x=7
7+1=8
Hence, the two integers are 7 and 8.
Paul.